If we consider the effect of taxes, then the degree of operating leverage can be written as: DOL = 1 + [FC × (1 – TC) – TC × D]/OCF Consider a project to supply Detroit with 28,000 tons of machine screws annually for automobile production. You will need an initial $6,100,000 investment in threading equipment to get the project started; the project will last for 6 years. The accounting department estimates that annual fixed costs will be $1,475,000 and that variable costs should be $280 per ton; accounting will depreciate the initial fixed asset investment straight-line to zero over the 6-year project life. It also estimates a salvage value of $850,000 after dismantling costs. The marketing department estimates that the automakers will let the contract at a selling price of $398 per ton. The engineering department estimates you will need an initial net working capital investment of $590,000. You require a return of 12 percent and face a tax rate of 23. a. What is the percentage change in OCF if the units sold changes to 29,000? (Do not round intermediate calculations and enter your answer as a percent rounded to 4 decimal places, e.g., 32.1616.) b. What is the DOL at the base-case level of output? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut aCopy Σ AutoSum Calibri ー E ゴWrap Text General ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B l u. Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard Font Alignment Number Styles Cells Edting BN388 BK BL BM BN BO BP BQ BR BS BT BU BV BW BX 388 389 390 391 392 393 394 395 396 397 398 399 400 401 402 403 404 405 406 4 EV PE EPS DIV ABNORMAL APV ROSS deberk MARGIN MONEY DETROIT DEPRECIATION 1016666.67 STEP 1 OCF ((sp-vc)(units) - fixed cost)(1-t) t(depreciation) OCF (398-280) (28000) 1475000) 1-0.23)(0.23) (1016666.67) OCF = 1642163.33 STEP 2 OCF ((sp-vc)(units) - fixed cost)(1-t) t(depreciation) OCF (398-280) (29000) 1475000) 1-0.23)(0.23) (1016666.67) OCF = 1733023.33 ANSI 96 CHANGE IN OCF- 5.5329% (BN396-BN392)/BN392 ANS 2 DOL1 (FIXED COST (1-t) DEPRECIATION(t))/OPERATING CASH FLOW 1(1475000(1-0.23) (1016666.67 0.23))/1642163.33 DOL- DOL = 1.5492 al tax . PB ROPİ te- , EVA MVA , LEVERAGED BUY SWAP-dstress . BRIGHAN . CAP GAIN erences: AL125 09:00